# Code-in-the-Loop Forensics: Agentic Tool Use for Image Forgery Detection

### v1

@@ 3.2.2 Cold Start
The training process for ForenAgent consists of two sequential stages, designed to progressively equip the MLLM with the capabilities to handle complex IFD tasks.
System Prompt Design: To steer the MLLM’s reasoning and code generation, ForenAgent uses a carefully engineered system prompt in addition to user queries. The prompt specifies how to access inputs, structure code, and return the final answer: (i) encourage executable code over free-form text; (ii) preload images/frames as image_clue_i (with resolution) so the model can reference them directly (e.g., cropping); (iii) standardize outputs via print(...) (text) and plt.show() (visuals); (iv) wrap every code block in <code>...</code> for reliable parsing; (v) place the final class token inside <answer>...</answer> for consistent evaluation. This design yields parsable, executable code with minimal runtime errors. The full system prompt appears in the Supplementary material.
Correct Reasoning Trajectories: Built on FABench, we curate a long-horizon, multi-turn instruction-tuning set for IFD to inject domain reasoning and long-CoT skills into open-source MLLMs. For each sample, we provide the System Prompt, Question, and images to GPT-4.1 [openai2025gpt41] to obtain an authenticity judgment and a multi-step chain. We retain a response only if: (1) the predicted label is correct; (2) it contains executable Python wrapped in <code>...</code>; (3) for tampered cases, the answer explicitly names the forged object. The filtered subset, containing approximately 200k agent–interaction question–answer pairs, is used for supervised Cold-Start tuning.
@@ 3.2.3 Reinforcement Fine-Tuning
Following DeepEyes [zheng2025deepeyes] and related work, we study how MLLMs acquire tool-calling and reasoning without supervised labels, using pure RL for self-improvement. End-to-end, outcome-rewarded RL jointly optimizes textual CoT and action planning over full trajectories. The agent interacts for multiple turns until producing an answer or exhausting the tool-call budget. States interleave text tokens $X$ and image tokens $I$; all observation tokens are inputs only and do not contribute to the loss.
Group Relative Policy Optimization (GRPO): With GRPO [guo2025deepseek], we sample a small candidate set per input, score them within-group, and update the policy without a critic using clipped importance weights plus a KL penalty to a reference, stabilizing training and leveraging preferences from model- or human-generated answers.
Reward Modeling: In addition to the correctness reward $R_{\text{acc}}(\tau)$ and the format reward $R_{\text{format}}(\tau)$ that ensures the use of valid <code> and <answer> tags, we introduce a tool usage reward $R_{\text{tool}}$ to evaluate how effectively the model applies external tools. The tools are categorized into Basic Image Processing $\mathcal{T}{\text{basic}}$ and Low-Level Forensics $\mathcal{T}{\text{low}}$. The reward $R_{\text{tool}}(\tau)$ integrates four components to assess the logical and context-aware use of these tools.
@@
By incorporating these verifiable reward components into the reinforcement learning process, ForenAgent achieves more interpretable and systematic reasoning for image forensic detection, effectively learning to leverage both basic image processing and low-level forensic analysis tools in a coherent investigative workflow.
4 Experiments
4.1 Experimental Setup
Baselines. We compare against 18 competitive baselines in three groups: (1) Closed-source MLLMs: Gemini2.5-flash [team2023gemini], Gemini2.5-Pro [team2023gemini], GPT-4o [achiam2023gpt], GPT-o3-mini [jaech2024openai], GPT-4.1 [openai2025gpt41], GPT-5 [openai2025gpt5]. (2) Large-scale MLLMs (zero-shot, no finetuning): InternVL3-78B [zhu2025internvl3], Qwen2.5-VL-72B [Qwen2.5-VL], QVQ-72B-preview [qvq-72b-preview], InternVL2.5-78B-MPO [wang2024mpo], Qwen3-VL-30B [Qwen2.5-VL]. (3) Trained baselines (same image training set as ForenAgent): we train Qwen2.5-VL-7B [Qwen2.5-VL] and Qwen3-VL-8B [Qwen2.5-VL] as MLLM comparison models. We also include advanced detectors Gram-Net [DBLP:conf/cvpr/LiuQT20], UnivFD [DBLP:conf/cvpr/OjhaLL23], LGrad [DBLP:conf/cvpr/Tan0WGW23], LNP [DBLP:journals/corr/abs-2311-00962], and SIDA (We use the 13B-parameter variant.) [huang2025sidasocialmediaimage].
Implementation Details. All experiments are conducted with PyTorch on eight NVIDIA Tesla H200 GPUs; we adopt Qwen2.5-VL-7B as the base MLLM, perform full-parameter finetuning in the Cold-Start stage with a learning rate of 1e-5 for two epochs using AdamW and a cosine-annealing scheduler with a maximum context length of 100k tokens, and then run RFT with GRPO on Qwen2.5-VL-7B for 80 iterations, sampling 256 prompts per batch with eight rollouts per prompt and at most seven tool calls and capping the response length at 20,480 tokens.
Evaluation Metrics. Following prior work [huang2025sidasocialmediaimage], we evaluate detection at the image level using Accuracy and F1. More details are provided in the Supplementary Material.
Table 1: Comparison of ForenAgent with other state-of-the-art methods on the FABench dataset.
@@ 4.1 Experimental Setup
GPT-o3-mini [jaech2024openai] 46.6 38.1 96.0 54.5 72.5 83.5 21.7 34.5 69.9 72.4 15.7 25.8 44.5 38.3
@@
Qwen3-VL-30B [Qwen2.5-VL] 55.8 42.4 90.6 57.7 65.4 45.5 19.7 27.5 71.7 67.7 29.0 40.6 46.4 41.9
Qwen2.5-VL-7B [Qwen2.5-VL] 86.2 80.2 78.0 79.1 86.1 79.5 78.4 78.9 87.1 79.5 82.7 81.1 79.7 79.7
Qwen3-VL-8B [Qwen2.5-VL] 80.1 66.2 81.9 73.2 86.9 88.7 69.6 78.0 85.2 78.4 76.9 77.6 76.1 76.3
@@
[…]
@@
[…]
@@ 4.2 Detection Evaluation
As shown in Table 1, we report performance across all test splits, and ForenAgent achieves the highest overall accuracy and F1-score, while also outperforming all baselines on both the synthetic and tampered categories. Notably, zero-shot performance from both closed-source and large-scale open-source MLLMs remains weak: these models consistently misclassify tampered and synthetic images as authentic, highlighting the insufficiency of IFD-relevant knowledge in current MLLM pretraining corpora. After supervised training, Qwen2.5-VL-7B outperforms Qwen3-VL-8B, supporting our choice of Qwen2.5-VL as the backbone. UnivFD achieves the best accuracy on the authentic class, suggesting strong capability in identifying unmanipulated images, but it struggles to distinguish between tampered and synthetic cases. Overall, these results underline the robustness of ForenAgent across manipulation types and domains, demonstrating its potential as a general-purpose and high-performance solution for real-world IFD.
@@ 4.3 Generalization
To assess generalization, we further evaluate ForenAgent on the SIDA-Test dataset in Table 2. Under identical training data, we compare ForenAgent with Gram-Net, UnivFD, LGrad, LNP, and SIDA. ForenAgent achieves the highest scores, demonstrating its strong adaptive capacity. We further discuss in the Supplementary Material the limitations of current MLLMs in using bounding boxes for forgery localization and outline the optimal pipeline for adapting ForenAgent to pixel-level localization.
@@ 4.4 Ablation Study
Methods Accuracy F1-score
w/o Cold Start 79.8 78.7
w/o RFT 81.9 81.7
w/o Tool Reward 83.6 83.5
[…]
Figure 4: The complete evidence chain by which ForenAgent correctly identifies a synthetic image.
@@ 4.5 Visualization
Figure 4 illustrates a successful synthetic-image case where ForenAgent constructs a coherent evidence chain that mirrors human reasoning, progressing through global perception, local focusing, iterative probing, and holistic adjudication to deliver both accurate detection and a convincing explanation. It first applies DCT-based global screening and flags initial suspicion, then conducts local focusing and iteratively probes with JPEG Ghost and Bayar Constrained Convolution, and finally integrates all cues into a well-founded decision. This example shows that ForenAgent not only produces correct labels but also assembles a logically sound, tool-driven forensic rationale, charting a promising path toward general-purpose IFD.
Figure 5: Distribution of tool usage frequencies.
@@ 4.6 Tool Usage Analysis
Figure 5 analyzes tool usage frequencies on the FABench test sets. For synthetic images, ForenAgent typically converges around 3 tool calls, whereas tampered images require around 4 calls due to their higher structural complexity and the need to localize manipulated regions. Figure 6 summarizes the usage distribution across low-level forensic tools: SRM, FFT, and JPEG Ghost are used most frequently, while PRNU and High-pass Filter appear less often. This pattern suggests that ForenAgent learns an adaptive tool-selection policy conditioned on image characteristics, rather than relying on mechanical tool enumeration. Notably, many classical low-level forensic tools are revived and integrated into our pipeline, offering a new perspective on combining traditional IFD techniques with modern MLLMs and potentially inspiring future hybrid-agent designs.
Figure 6: Low-level forensic tool usage frequency distribution across the FABench dataset.
@@ 5 Conclusion
In this paper, we introduced ForenAgent, an interactive multi-round framework that enables MLLMs to autonomously construct and iteratively refine Python-based low-level tools for image forgery detection. We abstract and generalize key low-level Python tools in IFD, forming a 12-tool forensic toolbox for future community extension. Through a two-stage training pipeline of Cold Start and Reinforcement Fine-Tuning, ForenAgent learns a dynamic reasoning process from global perception to holistic adjudication. Experimental results on FABench and SIDA-Test demonstrate its superior interpretability, robustness, and reflective tool-use capability across diverse forgery scenarios. Our work marks an important first step toward building intelligent agent systems for image forensics.

### v2

@@ 3.2.2 Cold Start
The training process for ForenAgent consists of two sequential stages, designed to progressively equip the MLLM with the capabilities to handle complex image forgery detection tasks.
System Prompt Design: To steer the MLLM’s reasoning, code generation, and low-level tool usage, ForenAgent employs a carefully designed system prompt in addition to user queries. The prompt defines how to access inputs, invoke tools, structure code, and return results: (i) prefer executable code or tool calls over free-form text; (ii) preload images or frames as image_clue_i (with resolution metadata) to enable direct referencing for operations such as cropping; (iii) standardize outputs via print(...) for text and plt.show() for visualizations; (iv) wrap each code block with <code>...</code> for reliable parsing, and each tool invocation with <tool_call>...</tool_call> to enable structured execution; (v) place the final class token inside <answer>...</answer> for consistent evaluation. This design produces parsable and executable outputs with reduced runtime errors. The full system prompt is provided in the Appendix.
Correct Reasoning Trajectories: Built on FABench, we curate a long-horizon, multi-turn instruction-tuning set for IFD to inject domain reasoning and long-CoT skills into open-source MLLMs. For each sample, we provide the System Prompt, question, and images to GPT-4.1 to obtain an authenticity judgment and a multi-step chain. We retain a response only if: (1) the predicted label is correct; (2) It includes a code sandbox that enforces the enclosure of executable Python within <code>...</code> and tool invocations within <tool_call>...</tool_call>; (3) for tampered cases, the answer explicitly names the forged object. The filtered subset, containing approximately 200k agent–interaction question–answer pairs, is used for supervised Cold-Start tuning.
@@ 3.2.3 Reinforcement Fine-Tuning
In this section, we investigate how MLLMs acquire tool-calling and reasoning capabilities without the need for supervised labels, leveraging pure RL for continuous self-improvement. End-to-end, outcome-rewarded RL jointly optimizes textual CoT and action planning over full trajectories. The agent interacts for multiple turns until producing an answer or exhausting the tool-call budget. States interleave text tokens $X$ and image tokens $I$; all observation tokens are inputs only and do not contribute to the loss.
Group Relative Policy Optimization (GRPO): With GRPO [guo2025deepseek], we sample a candidate set for each input, compute relative rewards through within-group normalization, and update the policy without a separate critic network. By employing clipped importance weights alongside a KL divergence penalty relative to a reference model, we stabilize the training process while effectively distilling preference signals from both model-generated and human-annotated answers.
Reward Modeling: In addition to the correctness reward $R_{\text{acc}}(\tau)$ and the format reward $R_{\text{format}}(\tau)$ that ensures the use of valid <code>,<tool_call> and <answer> tags, we introduce a tool usage reward $R_{\text{tool}}$ to evaluate how effectively the model applies external tools. The tools are categorized into Basic Image Processing $\mathcal{T}_{\text{basic}}$ and Low-Level Forensics $\mathcal{T}_{\text{low}}$. The reward $R_{\text{tool}}(\tau)$ integrates four components to assess the logical use of these tools.
@@
By incorporating these verifiable reward components into the reinforcement learning process, ForenAgent achieves more interpretable and systematic reasoning for IFD, effectively learning to leverage both basic image processing and low-level forensic analysis tools in a coherent investigative workflow.
Table 1: Comparison of ForenAgent with state-of-the-art methods on FABench, reporting per-class and overall performance across three categories.
@@
GPT-o3-mini [openai2025o3mini] 46.6 38.1 96.0 54.5 72.5 83.5 21.7 34.5 69.9 72.4 15.7 25.8 44.5 38.3
@@
Qwen3-VL-30B [Qwen3-VL] 55.8 42.4 90.6 57.7 65.4 45.5 19.7 27.5 71.7 67.7 29.0 40.6 46.4 41.9
GPT-4.1 (with Tools) [openai2025gpt41] 77.8 65.5 70.4 67.9 75.1 64.1 57.9 60.8 76.4 64.3 65.7 65.0 64.7 64.6
Qwen2.5-VL-72B (with Tools) [Qwen2.5-VL] 74.2 61.5 60.1 60.8 71.2 56.7 57.4 57.1 69.5 54.2 54.7 54.4 57.4 57.4
Gemini2.5-Pro (with Tools) [team2023gemini] 71.3 57.0 57.1 57.1 69.2 54.1 50.9 52.4 68.3 52.3 55.3 53.8 54.4 54.4
Qwen2.5-VL-7B [Qwen2.5-VL] 86.2 80.2 78.0 79.1 86.1 79.5 78.4 78.9 87.1 79.5 82.7 81.1 79.7 79.9
Qwen3-VL-8B [Qwen3-VL] 80.1 66.2 81.9 73.2 86.9 88.7 69.6 78.0 85.2 78.4 76.9 77.6 76.1 76.3
@@
ForenAgent 94.0 87.9 95.3 91.4 93.6 91.3 89.4 90.3 91.3 89.5 83.7 86.5 89.5 89.4
4 Experiments
4.1 Experimental Setup
Baselines. We compare ForenAgent against 24 baselines across four groups: (1) closed-source MLLMs evaluated zero-shot without fine-tuning, (2) large-scale open-source MLLMs evaluated zero-shot without fine-tuning, (3) tool-augmented MLLMs equipped with our toolbox, and (4) trained baselines, including MLLMs trained on the same image set as ForenAgent and representative state-of-the-art forgery detectors. Detailed baseline descriptions and configurations are provided in the Supplementary Materials. Notably, Our evaluation adopts a more challenging three-class protocol; most conventional detectors are binary by design and thus cannot be directly applied without substantial modification.
Implementation Details. All experiments are conducted with PyTorch on eight NVIDIA Tesla H200 GPUs; we adopt Qwen2.5-VL-7B as the base MLLM, perform full-parameter finetuning in the Cold-Start stage with a learning rate of 1e-5 for two epochs using AdamW and a cosine-annealing scheduler with a maximum context length of 100k tokens, and then run RFT with GRPO on Qwen2.5-VL-7B for 80 iterations, sampling 256 prompts per batch with eight rollouts per prompt and at most seven tool calls and capping the response length at 20,480 tokens. More details are provided in the Supplementary Materials.
Evaluation Metrics. Following prior work [huang2025sidasocialmediaimage], we evaluate detection at the image level using Accuracy and F1.
@@ 4.1 Experimental Setup
ForenAgent 83.1 82.9
Methods Accuracy F1-score
w/o Cold Start 78.7 76.9
w/o RFT 81.4 81.3
w/o Tool Reward 83.8 82.9
w/o global 86.4 86.3
w/o logic 84.4 84.3
w/o crop 87.8 86.5
w/o coh 88.0 87.9
with InternVL3-8B 79.8 79.9
with LLaVA-7B 75.8 75.5
ForenAgent 89.5 89.4
@@ 4.2 Detection Evaluation
As shown in Table 1, ForenAgent achieves state-of-the-art performance across all test splits, consistently outperforming baselines in both synthetic and tampered categories. In contrast, existing closed and open-source MLLMs struggle in zero-shot settings by frequently misclassifying manipulated images as authentic. This highlights a critical deficit of IFD-specific knowledge in their pretraining corpora. Notably, equipping models such as Gemini2.5-Pro and GPT-4.1 with specialized forensic tools consistently improves performance, underscoring the potential of tool-augmented IFDL. After supervised training, Qwen2.5-VL-7B outperforms Qwen3-VL-8B, supporting our choice of Qwen2.5-VL as the backbone. UnivFD achieves the best accuracy on the authentic class, suggesting strong capability in identifying unmanipulated images, but it struggles to distinguish between tampered and synthetic cases. Overall, these results underline the robustness of ForenAgent across manipulation types and domains, demonstrating its potential as a general-purpose and high-performance solution for real-world IFD.
@@ 4.3 Generalization
To assess generalization, we further evaluate ForenAgent on the SIDA-Test dataset in Table 3. Under identical training data, we compare ForenAgent with other methods. ForenAgent achieves the highest scores, demonstrating its strong adaptive capacity. The Supplementary Materials further explore the limitations of current MLLMs in using bounding boxes for forgery localization and detail the optimal pipeline for adapting ForenAgent to pixel-level precision.
@@ 4.4 Ablation Study
We also investigate the impact of specific reward components and backbone architectures. Systematic exclusion of the four reward functions: Global Forensic Priority ($R_{\text{global}}$), Tool Logic ($R_{\text{logic}}$), Crop Sensitivity ($R_{\text{crop}}$), and Reasoning Coherence ($R_{\text{coh}}$). Notably, the w/o logic variant exhibits the most significant drop in both accuracy and F1-score, suggesting that maintaining syntactically and logically correct tool usage is the foundation of agent efficacy. The degradation in w/o global and w/o coh further validates our "locate-then-investigate" reasoning chain. Moreover, replacing the default architecture with InternVL (with InternVL3-8B) or LLaVA (with LLaVA-7B) results in substantial performance loss, which confirms that our selected backbone is better suited for the IFD task. Collectively, these findings underscore the critical synergy between our staged training, structured reward design, and backbone selection.
Figure 5: The evidence chain by which ForenAgent correctly identifies a synthetic image.
Figure 6: Distribution of tool invocation frequencies for ForenAgent across FABench.
@@ 4.5 Visualization
Figure 5 illustrates a successful synthetic-image case where ForenAgent constructs a coherent evidence chain mirroring human reasoning. The process progresses through global perception, local focusing, iterative probing, and holistic adjudication to deliver accurate detection and a convincing explanation. It first applies the Bayar filter for global scanning to identify initial suspicion based on global texture uniformity and lack of photographic noise. The agent then conducts local focusing on a representative region and iteratively probes with DCT high-pass filtering and PRNU residual analysis. Finally, the agent synthesizes multi-scale forensic cues, concluding that the hyper-realistic smooth rendering and lack of random noise point to a fully synthetic origin.
@@ 4.6 Tool Usage Analysis
Figure 7: Quantitative distribution of low-level forensic tool usage frequencies across the comprehensive FABench dataset.
Figure 6 analyzes tool usage frequencies on the FABench test sets. For synthetic images, ForenAgent typically converges within about $3$ tool calls, whereas tampered images require around $4$ calls due to higher structural complexity and the need to localize manipulated regions. This suggests that ForenAgent dynamically adapts its reasoning workflow to task difficulty, substantially improving efficiency. Figure 7 summarizes the usage distribution across low-level forensic tools: SRM, FFT, and JPEG Ghost are used most frequently, while Laplacian High-Pass appears less often. This pattern suggests that ForenAgent learns an adaptive tool-selection policy conditioned on image characteristics, rather than relying on mechanical tool enumeration. Notably, many classical low-level forensic tools are revived and integrated into our pipeline, offering a new perspective on combining traditional image forgery detection techniques with modern MLLMs and potentially inspiring future hybrid-agent designs. Detailed evaluations of computational efficiency and robustness are discussed in the Supplementary Materials.
@@ 5 Conclusion
In this paper, we introduced ForenAgent, an interactive multi-round framework that enables MLLMs to autonomously invoke and iteratively refine Python-based low-level tools for IFD. We abstract and generalize key low-level tools in IFD, forming a 12-tool forensic toolbox for future community extension. Through a two-stage training pipeline of Cold Start and Reinforcement Fine-Tuning, ForenAgent learns a dynamic reasoning process from global perception to holistic adjudication. Experimental results on FABench and SIDA-Test demonstrate its superior interpretability, robustness, and reflective tool-use capability across diverse forgery scenarios. Our work marks an important first step toward building intelligent agent systems for image forensics.